OpenStudy (hailey):

The midpoint of a line segment with end points as (2, 3) and (4, y2) is (3, 2). What is the value of y2?

OpenStudy (hailey):

?

OpenStudy (anonymous):

(3,2) = ( (4+2)/2) , ( (y2+ 3)/2) 2 = (y2 + 3)/2 4 = y2 +3 y2= 1

OpenStudy (hailey):

thanks !!